# A Tutte-type canonical decomposition of 4-connected graphs

### v1

@@ Abstract.
As a corollary, we obtain a new Tutte-type canonical decomposition of 3-connected graphs into parts that are either quasi-4-connected, generalised wheels or thickened $K_{3,m}$’s. This decomposition is similar yet different from the tri-separation decomposition [Tridecomp].
As an application of the decomposition for 4-connectivity, we obtain a new theorem characterising all 4-connected vertex-transitive finite graphs as quasi-5-connected, the $K_{4}$-expansion of a quasi-5-connected graph, or on a short explicit list of graphs.
@@
[Uncaptioned image] Funded by DFG, project number 546892829.
@@ 1. Introduction
Is there for every $k\in\mathbb{N}$ an explicit canonical way to decompose every $k$-connected graph along $k$-separators into smaller pieces that are $(k+1)$-connected or ‘basic’?
@@
1.1 has a trivial solution for $k=0$, and for $k=1$ a positive answer is provided by the block-cutvertex-decomposition. Already for $k=2$, the solution is a classical theorem of Tutte [TutteCon] from 1966. It was not known how to extend Tutte’s approach to $k=3$ until recently, when Carmesin and Kurkofka discovered a new notion of 3-separators which allowed them to find a Tutte-type canonical decomposition of 3-connected graphs [Tridecomp]. However, examples show that the obvious extension of their solution to $k=4$ fails. In particular, it was not clear whether there even exists a notion of 4-separators that allows for such an extension.
In this paper, we introduce a new type of 4-separator and, as our main result, use it to prove a Tutte-type canonical decomposition theorem for 4-connectivity. As a corollary, we obtain a new Tutte-type canonical decomposition for 3-connectivity that provides an alternative perspective to [Tridecomp]. We also obtain a new theorem characterising vertex-transitive 4-connected finite graphs.
Background. To solve 1.1 for $k=0$ and $k=1$, one simply cuts along all $k$-separators of the graph. But already for $k=2$, this approach fails. Indeed, different 2-separators can cross in that they cut through each other, as happens in a cycle. Then if we cut at one of the 2-separators, we loose the other. Tutte realised that the solution is to cut the graph at all 2-separators that are totally-nested in that they are not crossed by any 2-separators, and indeed no conflicts arise when cutting in this way. The resulting pieces of Tutte’s decomposition are 3-connected, cycles or $K_{2}$’s. All three decompositions for $k\leqslant 3$ can be described in terms of tree-decompositions. For instance, Tutte’s selection of 2-separators gives a recipe for a tree-decomposition with all adhesion-sets of size $2$ and all whose torsos [^1] A torso of a tree-decomposition is obtained from a bag by making each of its adhesion-sets into a clique. are 3-connected, cycles or $K_{2}$’s.
@@
There are $k$-connected graphs which just slightly fail to be $(k+1)$-connected, but all whose $k$-separators are crossed. For example, if $G$ is $k$-connected and the neighbourhoods of its vertices are precisely the $k$-separators of $G$, then the $k$-separators around neighbouring vertices cross, but $G$ fails to be $(k+1)$-connected. Such $G$ exist for all $k\geqslant 3$: take a circular saw as in Figure 1.1, for instance.
Figure 1.1. A circular saw has vertex-set $\mathbb{Z}_{n}\times\mathbb{Z}_{2}$ with $n\gg k$ and every vertex $(v,0)$ sends edges to the $k$ vertices $(v,1),\ldots,(v+k-1,1)$. It is $k$-connected, and when $k\geqslant 3$ its neighbourhoods of vertices are precisely its $k$-separators (Lemma A.2).
@@
We consider the $k$-connected graph $G$ depicted in Figure 1.2 if $k$ is odd and in Figure 1.3 if $k$ is even. Every $k$-separator of $G$ is formed from a red set, a non-adjacent blue set, and if $k$ is even also the apex-vertex in the centre. Hence every $k$-separator of $G$ is crossed by another $k$-separator. However, $G$ is far from $(k+1)$-connected, and in no way ‘basic’: every $42k$-clique can be replaced with an arbitrarily large $(k+1)$-connected graph, so $G$ represents a whole zoo of examples.
@@
Decomposing 4-connected and 3-connected graphs. We found a way to overcome the two challenges for $k=4$ and $k=3$ with a twofold approach, which we first explain for $k=4$.
Ad 1. We relax the notion of $5$-connected to quasi-5-connected, an idea that we learned about from Grohe’s work [grohe2016quasi]. A graph $G$ is quasi-5-connected if it is 4-connected and every 4-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|=1$ or $|B\smallsetminus A|=1$.
Ad 2. We introduce the new notion of a tetra-separation, which we use instead of 4-separators. There are two key differences. On the one hand, tetra-separations may use edges in addition to vertices to cut up the graph, an idea that we learned about from the work of Carmesin and Kurkofka [Tridecomp] (we will discuss the relation later). On the other hand, every vertex used by a tetra-separation to cut up the graph into two sides must have at least two neighbours in each side, not counting neighbours that are also used by the tetra-separation. This degree-condition is new.
Recall that a mixed-separation of a graph $G$ is a pair $(A,B)$ with $A\cup B=V(G)$ and $A\smallsetminus B\neq\emptyset\neq B\smallsetminus A$. The separator of $(A,B)$ is the set of all vertices in $A\cap B$ and all edges in $G$ with one end in $A\smallsetminus B$ and the other in $B\smallsetminus A$. A tetra-separation of $G$ is a mixed-separation $(A,B)$ of $G$ whose separator has size four and which satisfies the following two conditions:
  1. degree-condition:
    every vertex in the separator has $\geqslant 2$ neighbours in both $A\smallsetminus B$ and $B\smallsetminus A$;
  2. matching-condition:
    the edges in the separator form a matching.
Two tetra-separations $(A,B)$ and $(C,D)$ are nested if, after possibly exchanging the name $A$ with $B$ or the name $C$ with $D$, we have $A\subseteq C$ and $B\supseteq D$. A tetra-separation of $G$ is totally-nested if it is nested with every tetra-separation of $G$. Figure 1.4 shows the totally-nested tetra-separations of an example-graph similar to the one shown in Figure 1.3.
@@
While the decompositions for $k$-connectivity for $k\leqslant 2$ such as the Tutte-decomposition can be described as tree-decompositions, the decomposition of 4-connected graphs via tetra-separations requires a notion that sits in between the notions of tree-decomposition (which allows only vertices in adhesion-sets) and Wollan’s notion of tree-cut decomposition (which allows only edges in adhesion-sets) [WollanExcludedImmersion]. To this end, the notion of mixed-tree-decomposition has been introduced which allows both vertices and edges in adhesion-sets [Tridecomp]. The set of totally-nested tetra-separations of a 4-connected graph $G$ canonically corresponds to a unique mixed-tree-decomposition of $G$. See Section 2.2 for details.
In order to state our decomposition-result for 4-connected graphs, we need the following definitions to describe the ‘basic’ parts of the decomposition – those that are not quasi-5-connected. Recall that graph-decompositions generalise tree-decompositions in that the decomposition-tree may take the form of an arbitrary graph, such as a cycle, in which case we speak of a cycle-decomposition [canonicalGraphDec]. Let $X$ be a graph. We call $X$ a generalised double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that $X-u-v$ has a cycle-decomposition of adhesion 1 into $K_{2}$’s and triangles (Figure 1.5). We call $X$ a cycle of $X_{1}$-torsos, …, $X_{k}$-torsos where the $X_{i}$ are graphs, if $X$ is 4-connected and has a cycle-decomposition with all adhesion-sets of size 2 such that every torso is isomorphic to some $X_{i}$ (Figure 1.6). Recall that the only 3-connected graphs on $\leqslant 5$ vertices are $K_{4}$, the $4$-wheel, $K_{5}$ minus an edge, and $K_{5}$. A thickened $K_{4,m}$ is obtained from a $K_{4,m}$ by turning the left side into a $K_{4}$. A sprinkled $K_{4,m}$ is obtained from a $K_{4,m}$ by adding any number of edges to the left side.
@@
Figure 1.6. A cycle of triangle-torsos, $K_{4}$-torsos and $K_{5}$-torsos
@@
Both $N(G)$ and $\mathcal{T}(G)$ in Theorem 1 are canonical: if $\varphi\colon G\to G^{\prime}$ is a graph-isomorphism, then $N(\varphi(G))=\varphi(N(G))$ and $\mathcal{T}(\varphi(G))=\varphi(\mathcal{T}(G))$.
@@
Theorem 1 includes infinite graphs $G$ with no additional restrictions whatsoever. More generally, all graphs in this paper are allowed to be infinite, unless stated otherwise.
Theorem 1 implies a new Tutte-type canonical decomposition for 3-connectivity. A strict tri-separation is defined like a tetra-separation, except that the separator is required to have size three instead of four. A strict tri-separation of a graph $G$ is totally-nested if it is nested with every strict tri-separation of $G$. A generalised wheel with centre $v$ is a 3-connected graph $X$ with a vertex $v$ such that $X-v$ has a cycle-decomposition with all adhesion-sets of size 1 and into $K_{2}$’s and triangles.
@@
Both Remarks 1.2 and 1.3 also apply to 2.
Vertex-transitive graphs. It is well known that all vertex-transitive finite connected graphs are either 3-connected, cycles, or $K_{2}$’s [infiniteSPQR]. This fact has then been refined in [Tridecomp] by analysing the 3-connected outcome further; see [Tridecomp, Corollary 2]. However, [Tridecomp] leaves it open to gain new insights into vertex-transitive 4-connectivity. This is where we apply Theorem 1: to obtain a characterisation of vertex-transitive 4-connectivity. Recall that a graph $G$ is arc-transitive if for every two ordered pairs $(u_{i},v_{i})$ with $u_{i}v_{i}\in E(G)$ for $i=1,2$ there is an automorphism $\varphi$ of $G$ such that $\varphi(u_{1})=u_{2}$ and $\varphi(v_{1})=v_{2}$. The $K_{4}$-expansion of a 4-regular graph $G$ is obtained from $G$ by expanding every vertex to a $K_{4}$, and adding a perfect matching such that two $K_{4}$’s are joined by a matching-edge if and only if their original vertices in $G$ are joined by a $G$-edge. For other definitions related to 2, see Section 12.
@@
mainresultmaintransitive A 4-connected finite graph $G$ is vertex-transitive if and only if $G$ either is
@@
    a quasi-5-connected vertex-transitive graph,
@@
    the $K_{4}$-expansion of a quasi-5-connected arc-transitive graph,
@@
    a cycle of $K_{4}$-bags of length $\geqslant 4$ (Figure 1.7),
@@
    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 4$ (Figure 1.8),
@@
    a cycle of triangle-bags of length $\geqslant 6$ (Figure 1.9),
or $G$ is one of the following graphs or a $K_{4}$-expansion thereof:
@@
    a cycle of $K_{2,2}$-bags of length $\geqslant 3$ (Figure 1.10),
@@
    the line-graph of the cube (Figure 1.11).
2 strengthens a classical theorem of Godsil and Royle from geometric group theory in a special case [Godsil, Theorem 3.4.2].
@@
A cycle of triangle-bags
Figure 1.10. A cycle of $K_{2,2}$-bags and its $K_{4}$-expansion
Figure 1.11. The line-graph of the cube and its $K_{4}$-expansion
Applications that require canonicity and explicitness. There are applications for partial solutions to 1.1, and in particular for Theorem 1 and [Tridecomp], that absolutely require canonicity and explicitness, often both. Examples of these applications include:

### v2

# A Tutte-type canonical decomposition of 3- and 4-connected graphs

@@ Abstract.
As a corollary, we obtain a new Tutte-type canonical decomposition of 3-connected graphs into parts that are either quasi-4-connected, generalised wheels or thickened $K_{3,m}$’s. This decomposition is similar yet different from the tri-separation decomposition [CK23].
As an application of the decomposition for 4-connectivity, we obtain a new theorem characterising all quasi-4-connected vertex-transitive finite graphs as essentially quasi-5-connected or on a short explicit list of graphs.
@@
[Uncaptioned image] funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 566118291; 546892829
funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 546892829
@@ 1. Introduction
Is there for every $k\in\mathbb{N}$ an explicit canonical way to decompose every $k$-connected graph along kk-‘separators’ into smaller pieces that are $(k+1)$-connected or ‘basic’?
@@
˜1.1 has a trivial solution for $k=0$, and for $k=1$ a positive answer is provided by the block-cutvertex-decomposition. Already for $k=2$, the solution is a classical theorem of Tutte [T66] from 1966. A non-canonical solution for $k=3$ was found by Grohe [G16] in 2016. Still, it was not known how to extend Tutte’s approach to $k=3$ until recently, when Carmesin and Kurkofka discovered a new notion of 3-separators which allowed them to find a Tutte-type canonical decomposition of 3-connected graphs [CK23]. However, examples show that the obvious extension of their solution to $k=4$ fails. In particular, it was not clear whether there even exists a notion of 4-separators that allows for such an extension.
In this paper, we introduce a new type of 4-separator and, as our main result, use it to prove a Tutte-type canonical decomposition theorem for 4-connectivity. As a corollary, we obtain a new Tutte-type canonical decomposition result for 3-connectivity that provides an alternative perspective to [CK23]. We also obtain a new theorem that characterises vertex-transitive quasi-4-connected finite graphs.
Background. To solve ˜1.1 for $k=0$ and $k=1$, one simply cuts along all $k$-separators of the graph. But already for $k=2$, this approach fails. Indeed, different 2-separators can cross in that they cut through each other, as happens in a cycle. Then if we cut at one of the 2-separators, we loose the other. Tutte realised that the solution is to cut the graph at all 2-separators that are totally-nested in that they are not crossed by any 2-separators, and indeed no conflicts arise when cutting in this way. The resulting pieces of Tutte’s decomposition are 3-connected, cycles or $K_{2}$’s. All decompositions for $k\leqslant 2$ can be described in terms of tree-decompositions. For instance, Tutte’s selection of 2-separators gives a recipe for a tree-decomposition with all adhesion-sets of size $2$ and all whose torsos [^1] A torso of a tree-decomposition is obtained from a bag by making each of its adhesion-sets into a clique. are 3-connected, cycles or $K_{2}$’s.
@@
There are $k$-connected graphs which just slightly fail to be $(k+1)$-connected, but all whose $k$-separators are crossed. For example, if $G$ is $k$-connected and the neighbourhoods of its vertices are precisely the $k$-separators of $G$, then the $k$-separators around neighbouring vertices cross, but $G$ fails to be $(k+1)$-connected. Such $G$ exist for all $k\geqslant 3$: take a circular saw as in Figure˜1.1, for instance. We believe that such graphs should not have to be decomposed any further at level $k$.
Figure $1.$ A circular saw has vertex-set $\mathbb{Z}_{n}\times\mathbb{Z}_{2}$ with $n\gg k$ and every vertex $(v,0)$ sends edges to the $k$ vertices $(v,1),\ldots,(v+k-1,1)$. It is $k$-connected, and when $k\geqslant 3$ its neighbourhoods of vertices are precisely its $k$-separators (Lemma˜14.2).
@@
We consider the $k$-connected graph $G$ depicted in Figure˜1.2 if $k$ is odd and in Figure˜1.3 if $k$ is even. Every $k$-separator of $G$ is formed from a red set, a non-adjacent blue set, and if $k$ is even also the apex-vertex in the centre. Hence every $k$-separator of $G$ is crossed by another $k$-separator. However, $G$ is far from $(k+1)$-connected, and in no way ‘basic’: every $42k$-clique can be replaced with an arbitrarily large $(k+1)$-connected graph, so $G$ represents a whole zoo of examples.
@@
Decomposing 3- and 4-connected graphs. We found a way to overcome the two challenges for $k=3$ and $k=4$ with a twofold approach, which we first explain for $k=4$.
Ad ˜1. We relax the notion of $5$-connected to quasi-5-connected, an idea that we learned about from Grohe’s work [G16]. A graph $G$ is quasi-5-connected if it is 4-connected and every 4-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|=1$ or $|B\smallsetminus A|=1$. We will not decompose quasi-5-connected graphs any further.
Ad ˜2. We introduce the new notion of a tetra-separation, which we use instead of 4-separators. There are two key differences. On the one hand, tetra-separations may use edges in addition to vertices to cut up the graph, an idea that we learned about from the work of Carmesin and Kurkofka [CK23] (we will discuss the relation later). On the other hand, every vertex used by a tetra-separation to cut up the graph into two sides must have at least two neighbours in each side, not counting neighbours that are also used by the tetra-separation. This degree-condition is new.
A mixed-separation of a graph $G$ is a pair $(A,B)$ with $A\cup B=V(G)$ and $A\smallsetminus B\neq\emptyset\neq B\smallsetminus A$. The separator of $(A,B)$ is the set of all vertices in $A\cap B$ and all edges in $G$ with one end in $A\smallsetminus B$ and the other in $B\smallsetminus A$. In a 4-connected graph, every mixed-separation has a separator of size $\geqslant 4$ (Lemma˜6.1). A tetra-separation of $G$ is a mixed-separation $(A,B)$ of $G$ whose separator has size four and which satisfies the following two conditions:
degree-condition  : every vertex in the separator has $\geqslant 2$ neighbours in both $A\smallsetminus B$ and $B\smallsetminus A$;
matching-condition  : the edges in the separator form a matching.
Two tetra-separations $(A,B)$ and $(C,D)$ are nested if, after possibly exchanging the name $A$ with $B$ or the name $C$ with $D$, we have $A\subseteq C$ and $B\supseteq D$. A tetra-separation of $G$ is totally-nested if it is nested with every tetra-separation of $G$. Figure˜1.4 shows the totally-nested tetra-separations of an example-graph similar to the one shown in Figure˜1.3.
@@
While the decompositions for $k$-connectivity for $k\leqslant 2$ such as the Tutte-decomposition can be described as tree-decompositions, the decomposition of 4-connected graphs via tetra-separations requires a notion that sits in between the notions of tree-decomposition (which allows only vertices in adhesion-sets) and Wollan’s notion of tree-cut decomposition (which allows only edges in adhesion-sets) [W15]. To this end, the notion of mixed-tree-decomposition has been introduced which allows both vertices and edges in adhesion-sets [CK23]. The set of totally-nested tetra-separations of a 4-connected graph $G$ canonically corresponds to a unique mixed-tree-decomposition of $G$. See Section˜2.2 for details.
In order to state our decomposition-result for 4-connected graphs, we need the following definitions to describe the ‘basic’ parts of the decomposition – those that are not quasi-5-connected. Graph-decompositions generalise tree-decompositions in that the decomposition-tree may take the form of an arbitrary graph, such as a cycle, in which case we speak of a cycle-decomposition [DJKK22]. Let $X$ be a graph. We call $X$ a generalised double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that XuvX-u-v has a cycle-decomposition of adhesion 1 into $K_{2}$’s and triangles (Figure˜1.5). We call $X$ a cycle of $X_{1}$-torsos, …, $X_{k}$-torsos, $Y_{1}$-bags, …, $Y_{\ell}$-bags where the $X_{i}$ and $Y_{j}$ are graphs, if $X$ is 4-connected and has a cycle-decomposition with all adhesion-sets of size 2 such that for every bag either its torso is isomorphic to some $X_{i}$ or the bag itself is isomorphic to some $Y_{j}$ (Figure˜1.6). Recall that the only 3-connected graphs on $\leqslant 5$ vertices are $K_{4}$, the $4$-wheel, $K_{5}$ minus an edge, and $K_{5}$. A thickened $K_{4,m}$ is obtained from a $K_{4,m}$ by turning the left side into a $K_{4}$. A sprinkled $K_{4,m}$ is obtained from a $K_{4,m}$ by adding any number of edges to the left side.
@@
Figure 1.6. A cycle of triangle-bags, $K_{4}$-bags and $K_{5}$-bags
@@
Both $N(G)$ and $\mathcal{T}(G)$ in Theorem˜1 are canonical: if $\varphi\colon G\to G^{\prime}$ is a graph-isomorphism, then $N(\varphi(G))=\varphi(N(G))$ and $\mathcal{T}(\varphi(G))=\varphi(\mathcal{T}(G))$, and the automorphism group of every $G$ induces a group-action on $N(G)$ and on $\mathcal{T}(G)$.
@@
Theorem˜1 includes infinite graphs $G$ with no additional restrictions whatsoever. More generally, all graphs in this paper are allowed to be infinite, unless stated otherwise.
Theorem˜1 implies a new Tutte-type canonical decomposition for 3-connectivity. A strict tri-separation is defined like a tetra-separation, except that the separator is required to have size three instead of four. A strict tri-separation of a graph $G$ is totally-nested if it is nested with every strict tri-separation of $G$. A generalised wheel with centre $v$ is a 3-connected graph $X$ with a vertex $v$ such that XvX-v has a cycle-decomposition with all adhesion-sets of size 1 and into $K_{2}$’s and triangles.
@@
Both Remarks˜1.2 and 1.3 also apply to ˜2.
Comparison to the tri-separation decomposition. Carmesin and Kurkofka introduced the first Tutte-type canonical decomposition for 3-connected graphs [CK23]. They used tri-separations instead of strict tri-separations to obtain their decomposition. The main difference between tri-separations on the one hand, and strict tri-separations and tetra-separations on the other hand, is how their degree-conditions handle neighbours in the separator: tri-separations count them, but strict tri-separations and tetra-separations ignore them. This seemingly subtle difference has a surprisingly big impact: the verbatim extension of the tri-separation approach to 4-connectivity fails ˜2.
For 3-connected graphs, the strict tri-separation decomposition can be viewed as a refinement of the tri-separation decomposition, but the details are a bit technical; see Section˜15.
Algorithmic aspects. It is straightforward to come up with a polynomial-time algorithm that produces the decompositions in Theorem˜1 and ˜2 for all 4- and 3-connected graphs.
Korhonen [K24] recently showed that there is an algorithm that, given an $n$-vertex $m$-edge graph $G$ and an integer $k\geqslant 1$, returns a $k$-lean tree-decomposition of $G$ in time $k^{\mathcal{O}(k^{2})}n+\mathcal{O}(m)$. We believe that every 5-lean tree-decomposition of every 4-connected graph $G$ can be made into a tetra-separation decomposition of $G$ as in Theorem˜1 in linear time:
Conjecture 1.4.
There is an algorithm that, given a 4-connected $n$-vertex $m$-edge graph $G$, returns the tetra-separation decomposition of $G$ from Theorem˜1 in linear time $\mathcal{O}(n+m)$.
A positive answer would also provide a linear-time algorithm for the decomposition of 3-connected graphs in ˜2 by following the short proof of ˜2.
Total-nestedness makes parallel computing a feasible option for computing the tetra-separation decomposition even more efficiently. Indeed, all the partial solutions found in parallel, which come in the form of sets of totally-nested tetra-separations, can always be combined without conflict.
Combining the decompositions for different $k$. Every finite graph $G$ can be decomposed canonically by combining the partial solutions to ˜1.1 inductively, as follows.
  1. 1)
    First, we trivially decompose $G$ into its connected components.
  2. 2)
    Each connected component we further decompose with the block-cutvertex decomposition into bags that are 2-connected or $K_{2}$’s.
  3. 3)
    Each 2-connected bag we further decompose with the Tutte-decomposition into torsos that are 3-connected, cycles or $K_{2}$’s.
  4. 4)
    Each 3-connected torso we further decompose with ˜2 into torsos that are quasi-4-connected, generalised wheels or thickened $K_{3,m}$’s. Alternatively, we use [CK23].
[…]
Figure 1.7. A quasi-4-connected version of ˜2: all tetra-separations are crossed, yet the graph is far from 55-connected and each $K_{10}$ can be replaced with an arbitrary 5-connected graph
Let $H$ be an arbitrary graph. The following definition of $H^{\Delta}$ is supported by Figure˜1.8. Let $U$ be the set of all vertices in $H$ of degree three. We obtain $H^{s}$ from $H$ by subdividing every edge that joins two vertices in $U$. Then we obtain $H^{s\Delta}$ from $H^{s}$ by making each neighbourhood $N_{H^{s}}(u)$ around a vertex $u\in U$ into a triangle $\Delta(u)$. Finally, we let $H^{\Delta}:=H^{s\Delta}-U$. If $H$ is quasi-4-connected and has >6>6 vertices, then $H^{\Delta}$ is 4-connected by Lemma˜12.2. The operation $H\mapsto H^{\Delta}$ is canonical: every isomorphism $H_{1}\to H_{2}$ induces an isomorphism $H^{\Delta}_{1}\to H^{\Delta}_{2}$.
Figure 1.8. From $H$ (left) over $H^{s}$ and $H^{s\Delta}$ to $H^{\Delta}$ (right). Central graphs: vertices in $U$ are green, neighbours of $U$ are circled red, subdivision vertices are fully red, and the triangles $\Delta(u)$ are red.
[…]
  1. 5)
    Each quasi-4-connected torso $\tau$ with >6>6 vertices we make into a 4-connected graph $\tau^{\Delta}$.
    (Quasi-4-connected graphs on $\leqslant 6$ vertices are treated as ‘basic’ outcome.)
  2. 6)
    Finally, we decompose each 4-connected $\tau^{\Delta}$ with Theorem˜1.
The operation $\tau\mapsto\tau^{\Delta}$ harmonises well with the other decomposition steps. Indeed, we have followed all steps to prove Figure˜1.8 below.
Vertex-transitive graphs. Recall that a graph $G$ is arc-transitive if for every two ordered pairs $(u_{i},v_{i})$ with $u_{i}v_{i}\in E(G)$ for $i=1,2$ there is an automorphism $\varphi$ of $G$ such that $\varphi(u_{1})=u_{2}$ and $\varphi(v_{1})=v_{2}$. Let $H$ be an $r$-vertex graph and let $G$ be an $r$-regular graph. An $H$-expansion of $G$ is a graph obtained from $G$ by replacing the vertices $v$ of $G$ with pairwise disjoint copies $H_{v}$ of $H$ and by replacing every edge $e=uv$ of $G$ with an $H_{u}$$H_{v}$ edge $\hat{e}$ so that $\hat{e}_{1}$ and $\hat{e}_{2}$ share no ends for every two $e_{1}\neq e_{2}\in E(G)$. Thus, the edges $\hat{e}$ form a perfect matching of the $H$-expansion. It was mostly222[CK23, Corollary 2] is slightly weaker than the result we mention here. The slight strengthening happens in Proposition 13.51 in this paper. shown in [CK23] that a finite connected graph $G$ is vertex-transitive if and only if $G$ is either
  • $\cdot$
    quasi-4-connected and vertex-transitive,
  • $\cdot$
    the $K_{3}$-expansion of a quasi-4-connected 4-regular arc-transitive graph,
  • $\cdot$
    a cycle, $K_{2}$ or $K_{1}$.
[…]
A graph $G$ is 2-quasi-5-connected if $G$ is quasi-4-connected and every 4-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|\leqslant 2$ or $|B\smallsetminus A|\leqslant 2$. For other definitions related to Figure˜1.8, see Section˜13.
@@
mainresultmaintransitive A quasi-4-connected finite graph $G$ is vertex-transitive if and only if $G$ either is
@@
    a $(\geqslant 4)$-regular quasi-5-connected vertex-transitive graph,
@@
    a 3-regular 2-quasi-5-connected vertex-transitive graph,
@@
    a $K_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph, or a vertex-transitive $C_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph,
@@
    a cycle of $K_{4}$-bags of length $\geqslant 4$ (Figure˜1.9),
@@
    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 4$ (Figure˜1.10),
@@
    a cycle of triangle-bags of length $\geqslant 6$ (Figure˜1.12),
@@
    a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\geqslant 6$ (Figure˜1.11),
  8. (viii)
    a cycle of $C_{4}$-bags of length $\geqslant 4$ (Figure˜1.13) or $K_{3}\square K_{2}$,
  9. (ix)
    a cycle of $K_{2,2}$-bags of length $\geqslant 3$, the $K_{4}$-expansion thereof, or a vertex-transitive $C_{4}$-expansion thereof (Figure˜1.14), or
  10. (x)
    the line-graph of the cube, its $K_{4}$-expansion, or one of the three $C_{4}$-expansions of it depicted in Figure˜1.15.
@@
A cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos
Figure 1.12.
Cycles of triangle-bags
Figure 1.13.
Cycles of $C_{4}$-bags
Figure 1.14. A cycle of $K_{2,2}$-bags, its $K_{4}$-expansion, and two vertex-transitive $C_{4}$-expansions
Figure 1.15. The line-graph of the cube, its $K_{4}$-expansion, and its vertex-transitive $C_{4}$-expansions
When canonicity and an explicit description matter. There are applications for partial solutions to ˜1.1, and in particular for Theorem˜1 and [CK23], that require canonicity and explicitness, often both. Examples of these applications include: